# Gradient flows of $(K,N)$-convex functions with negative $N$

### v1

@@ 3.1  Definition of reparametrization maps
We denote by $\mathcal{C}$ the set of all the curves parametrized on some interval $(0,T)$, with $T\in(0,+\infty]$, and with image in $\mathcal{D}[f]$. We introduce the following two subsets of $\mathcal{C}$:
@@
$\mathcal{C}^{\prime\prime}_{N}\coloneqq\Big{\{}(y_{t})_{t\in(0,T)}\in\mathcal{C}^{\prime}\,:\,\int_{0}^{T\wedge 1}f_{N}(y_{t})\,\mathrm{d}t<\infty\Big{\}}.$ (40)
@@
where we set $T^{\prime}\coloneqq\lim_{t\to T^{-}}\alpha(t)$. Observe that $\alpha$ is continuous and almost everywhere differentiable with derivative
@@
In particular, $\alpha$ is strictly increasing and thus invertible. Let $\varphi\colon(0,T^{\prime})\to(0,T)$ be its inverse, for which we have $\lim_{s\to 0^{+}}\varphi(s)=0$, $\lim_{s\to T^{\prime-}}\varphi(s)=T$ and
$\varphi^{\prime}(s)=\frac{1}{\alpha^{\prime}(\varphi(s))}=-\frac{1}{N}f_{N}(y_{\varphi(s)})>0,\qquad s\in[0,T^{\prime}],$ (44)
so that also $\varphi$ is continuous and strictly increasing. Finally, we set
@@
where we set $T=\lim_{s\to T^{\prime}}\beta(s)$. Observe that, since $(z_{s})_{s\in(0,T^{\prime})}\in\mathcal{C}^{\prime\prime}_{N}$, $\beta$ is well defined. Moreover, it is continuous and everywhere differentiable with derivative
$\beta^{\prime}(s)=-\frac{f_{N}(y_{t})}{N}>0,\qquad\forall s\in(0,T^{\prime}),$ (48)
which ensures that it is strictly increasing and thus invertible. Let $\psi\colon(0,T)\to(0,T^{\prime})$ be its inverse, for which we have $\lim_{t\to 0^{+}}\psi(t)=0$, $\lim_{t\to T^{-}}\psi(t)=T^{\prime}$ and
$\psi^{\prime}(t)=\frac{1}{\beta^{\prime}(\psi(t))}=-\frac{N}{f_{N}(y_{\psi(t)})}>0,\qquad t\in(0,T),$ (49)
in particular $\psi$ is continuous and strictly increasing. Then, we set
@@
Given any $(y_{t})_{t\in(0,T)}\in\mathcal{C}^{\prime}$ and taking $(z_{s})_{s\in(0,T^{\prime})}=\mathcal{R}_{1}\big{(}(y_{t})_{t\in(0,T)}\big{)}$, we observe that the maps $\varphi$ and $\psi$ associated to these curves (respectively), are inverse to one another. As a consequence, we deduce that $\mathcal{R}_{2}\big{(}\mathcal{R}_{1}\big{(}(y_{t})_{t\in(0,T)}\big{)}\big{)}=(y_{t})_{t\in(0,T)}$ and thus that $\mathcal{R}_{1}\circ\mathcal{R}_{2}=\operatorname{id}_{\mathcal{C}^{\prime}}$. Analogously, we can prove that $\mathcal{R}_{2}\circ\mathcal{R}_{1}=\operatorname{id}_{\mathcal{C}^{\prime\prime}_{N}}$. ∎
@@ 3.2  Local characterization and self-improvement of $\operatorname{EVI}_{K,N}$ gradient curves
where the last inequality is a consequence of Lemma 3.5.
@@ 3. $(iii)$
    where $\tilde{M}\coloneqq-\frac{K}{N}e^{-M/N}$ and $M\coloneqq\max\{f(z_{r}),f(\tilde{z}_{r})\}$.
@@
$\mathfrak{s}_{K,N}(\theta/2)^{2}=-\frac{N}{K}\bigg{(}\mathfrak{c}_{K,N}(\theta)-1\bigg{)},$ (77)
@@
$\begin{split}\sigma_{K,N}^{(1-s)}(\mathsf{d})f_{N}(x_{0})&+\sigma_{K,N}^{(s)}(\mathsf{d})f_{N}(x_{1})\\ &\geq\frac{f_{N}(\gamma_{s}^{t})}{(e^{Kt}-1)}\bigg{[}\sigma_{K,N}^{(1-s)}(\mathsf{d})\big{(}e^{Kt}\mathfrak{c}_{K,N}(\mathsf{d}(\gamma_{s}^{t},x_{0}))-\mathfrak{c}_{K,N}((1-s)\mathsf{d})\big{)}\\ &\qquad\qquad\qquad\qquad+\sigma_{K,N}^{(s)}(\mathsf{d})\big{(}e^{Kt}\mathfrak{c}_{K,N}(\mathsf{d}(\gamma_{s}^{t},x_{0}))-\mathfrak{c}_{K,N}(s\mathsf{d})\big{)}\bigg{]},\end{split}$ (78)
for every $t$ sufficiently small. On the one hand, as a consequence of the trigonometric sum formulas, we observe that
$\sigma_{K,N}^{(1-s)}(\mathsf{d})\mathfrak{c}_{K,N}((1-s)\mathsf{d})+\sigma_{K,N}^{(s)}(\mathsf{d})\mathfrak{c}_{K,N}(s\mathsf{d})=1.$ (79)
On the other hand, using the trigonometric inequalities stated at the end of the proof of [EKS15, Thm. 2.23], we deduce that
$\sigma_{K,N}^{(1-s)}(\mathsf{d})\mathfrak{c}_{K,N}(\mathsf{d}(\gamma_{s}^{t},x_{0}))+\sigma_{K,N}^{(s)}(\mathsf{d})\mathfrak{c}_{K,N}(\mathsf{d}(\gamma_{s}^{t},x_{0}))\,\begin{cases}<1&\text{if }K<0,\\ >1&\text{if }K>0.\end{cases}$ (80)
@@ 5  $\operatorname{EVI}$ and $\operatorname{EDI}$
$f(y_{t})\leq\lim_{\epsilon\downarrow 0}f(y_{\epsilon})-\int_{0}^{t}\left[\tfrac{1}{2}|\dot{y}_{s}|^{2}+\tfrac{1}{2}|D^{-}f|^{2}(y_{s})\right]\,\mathrm{d}s\qquad\text{for every }t\in(0,T).$ (EDI)
@@ 3. $(c)$
    $-\tfrac{\,\mathrm{d}}{\,\mathrm{d}t}f(y_{t})\geq\tfrac{1}{2}|\dot{y}_{t}|^{2}+\tfrac{1}{2}|D^{-}f|^{2}(y_{t})\qquad\text{for }\mathscr{L}^{1}\text{-a.e. }t\in(0,T).$ (82)
@@
$\tfrac{\,\mathrm{d}}{\,\mathrm{d}t}f(y_{t})=-|D^{-}f|^{2}(y_{t})=-|\dot{y}_{t}|^{2}\qquad\text{for }\mathscr{L}^{1}\text{-a.e. }t\in(0,T).$ (83)
@@
$f(y_{t})=\lim_{\epsilon\downarrow 0}f(y_{\epsilon})-\int_{0}^{t}\left[\tfrac{1}{2}|\dot{y}_{s}|^{2}+\tfrac{1}{2}|D^{-}f|^{2}(y_{s})\right]\,\mathrm{d}s\qquad\text{for every }0<t<T.$ (EDE)
@@
Fix any $y\in\mathcal{D}[f]$ and $R>0$. We start by observing that
@@
$\begin{split}-\frac{N}{f_{N}(y)}\frac{f_{N}(\gamma_{t})-f_{N}(y)}{\mathsf{d}(y,\gamma_{t})}\leq&-\frac{N}{f_{N}(y)}\frac{\sigma_{K,N}^{(t)}(\mathsf{d})f_{N}(z)-(1-\sigma_{K,N}^{(1-t)}(\mathsf{d}))f_{N}(y)}{t\mathsf{d}}\\ \xrightarrow{t\to 0}&-\frac{N}{f_{N}(y)}\bigg{[}\frac{1}{\mathfrak{s}_{K,N}(\mathsf{d})}f_{N}(z)-\frac{\mathfrak{c}_{K,N}(\mathsf{d})}{\mathfrak{s}_{K,N}(\mathsf{d})}f_{N}(y)\bigg{]}\\ =&-\frac{N}{f_{N}(y)}\bigg{[}\frac{f_{N}(z)-f_{N}(y)}{\mathfrak{s}_{K,N}(\mathsf{d})}-\frac{1-\mathfrak{c}_{K,N}(\mathsf{d})}{\mathfrak{s}_{K,N}(\mathsf{d})}f_{N}(y)\bigg{]}\\ =&\frac{N}{\mathfrak{s}_{K,N}(\mathsf{d})}\left[1-\frac{f_{N}(z)}{f_{N}(y)}\right]-K\,\frac{\mathfrak{s}_{K,N}(\mathsf{d}/2)}{\mathfrak{c}_{K,N}(\mathsf{d}/2)}f_{N}(y),\end{split}$ (92)
@@
$\mathfrak{l}^{\Gamma}_{f}(y)\coloneqq\sup_{z\in\Gamma\setminus\{y\}}\frac{\big{(}f(z)-f(y)\big{)}^{-}}{\mathsf{d}(z,y)}.$
@@
$\mathfrak{l}^{\Gamma}_{f}(y)\leq|D^{-}f|(y)+|K|\max\bigg{\{}\frac{\mathfrak{s}_{K,N}(R/2)}{\mathfrak{c}_{K,N}(R/2)},\frac{\mathfrak{s}_{-K,N}(R/2)}{\mathfrak{c}_{-K,N}(R/2)}\bigg{\}},\qquad\forall y\in\Gamma,$
@@
Integrating (EDI) we deduce that $\left(y_{t}\right)_{t\in(0,T)}\in\mathcal{C}^{\prime}$. Consider the reparametrizations $\mathcal{R}_{1}\big{(}\left(y_{t}\right)_{t\in(0,T)}\big{)}$ and $\mathcal{R}_{1}\big{(}\left(y^{\prime}_{t}\right)_{t\in(0,T)}\big{)}$. Since $\mathcal{R}_{1}$ is injective on $\mathcal{C}^{\prime}$ by Proposition 3.1, it suffices to show that $\mathcal{R}_{1}\big{(}\left(y_{t}\right)_{t\in(0,T)}\big{)}=\mathcal{R}_{1}\big{(}\left(y^{\prime}_{t}\right)_{t\in(0,T)}\big{)}$.
@@
$-\tfrac{\,\mathrm{d}}{\,\mathrm{d}t}f_{N}(z_{t})$ $=-\frac{f_{N}(z_{t})}{N}\left[-\tfrac{\,\mathrm{d}}{\,\mathrm{d}t}f(z_{t})\right]=\varphi^{\prime}(t)\left[-\tfrac{\,\mathrm{d}}{\,\mathrm{d}t}f(z_{t})\right]$ (94)
@@
$-\tfrac{\,\mathrm{d}}{\,\mathrm{d}t}f(z_{t})=-\tfrac{\,\mathrm{d}}{\,\mathrm{d}t}f(y_{\varphi(t)})=\varphi^{\prime}(t)\tfrac{\,\mathrm{d}}{\,\mathrm{d}s}f(y_{s})\big{\rvert}_{s=\varphi(t)}\geq\varphi^{\prime}(t)\left[\tfrac{1}{2}\left\lvert\dot{y}_{\varphi(t)}\right\rvert^{2}+\tfrac{1}{2}\left\lvert D^{-}f\right\rvert^{2}(y_{\varphi(t)})\right].$ (95)
@@
$-\tfrac{\,\mathrm{d}}{\,\mathrm{d}t}f_{N}(z_{t})\geq\varphi^{\prime}(t)^{2}\left[\tfrac{1}{2}\left\lvert\dot{y}_{\varphi(t)}\right\rvert^{2}+\tfrac{1}{2}\left\lvert D^{-}f\right\rvert^{2}(y_{\varphi(t)})\right]=\tfrac{1}{2}\left\lvert\dot{z}_{t}\right\rvert^{2}+\tfrac{1}{2}\left\lvert D^{-}f_{N}\right\rvert^{2}(z_{t}).$ (96)

### v2

@@ 3.1  Definition of reparametrization maps
We denote by $\mathcal{C}$ the set of all the continuous curves parametrized on some interval $(0,T)$, with $T\in(0,+\infty]$, and with image in $\mathcal{D}[f]$. We introduce the following two subsets of $\mathcal{C}$:
@@
$\mathcal{C}^{\prime\prime}_{N}:=\Big{\{}(y_{t})_{t\in(0,T)}\in\mathcal{C}^{\prime}\,:\,\int_{0}^{T\wedge 1}f_{N}(y_{t})\,\mathrm{d}t<\infty\Big{\}}.$ (41)
@@
where we set $T^{\prime}\coloneqq\lim_{t\to T^{-}}\alpha(t)$. Observe that $\alpha$ is locally bi-Lipschitz and almost everywhere differentiable with derivative
@@
In particular, $\alpha$ is strictly increasing and thus invertible. Let $\varphi\colon(0,T^{\prime})\to(0,T)$ be its inverse, for which we have $\lim_{s\downarrow 0}\varphi(s)=0$, $\lim_{s\to T^{\prime-}}\varphi(s)=T$ and
$\varphi^{\prime}(s)=\frac{1}{\alpha^{\prime}(\varphi(s))}=-\frac{1}{N}f_{N}(y_{\varphi(s)})>0,\qquad\text{for a.e. }s\in[0,T^{\prime}],$ (45)
so that also $\varphi$ is locally bi-Lipschitz and strictly increasing. Finally, we set
@@
where we set $T=\lim_{s\to T^{\prime}}\beta(s)$. Observe that, since $(z_{s})_{s\in(0,T^{\prime})}\in\mathcal{C}^{\prime\prime}_{N}$, $\beta$ is well defined. Moreover, it is locally bi-Lipschitz and almost everywhere differentiable with derivative
$\beta^{\prime}(s)=-\frac{f_{N}(z_{s})}{N}>0,\qquad\forall s\in(0,T^{\prime}),$ (49)
which ensures that it is strictly increasing and thus invertible. Let $\psi\colon(0,T)\to(0,T^{\prime})$ be its inverse, for which we have $\lim_{t\downarrow 0}\psi(t)=0$, $\lim_{t\to T^{-}}\psi(t)=T^{\prime}$ and
$\psi^{\prime}(t)=\frac{1}{\beta^{\prime}(\psi(t))}=-\frac{N}{f_{N}(z_{\psi(t)})}>0,\qquad\text{for a.e. }t\in(0,T),$ (50)
in particular $\psi$ is locally bi-Lipschitz and strictly increasing. Then, we set
@@
Consider any $(y_{t})_{t\in(0,T)}\in\mathcal{C}^{\prime}$ and take $(z_{s})_{s\in(0,T^{\prime})}=\mathcal{R}_{1}\big{(}(y_{t})_{t\in(0,T)}\big{)}$. According to the discussion above, the maps $\varphi$ and $\psi$ associated to these curves (respectively), are locally bi-Lipschitz. Thus, their composition $\varphi\circ\psi:(0,T)\to(0,T)$ is itself locally bi-Lipschitz and, by (45) and (50), it satisfies
$(\varphi\circ\psi)^{\prime}(t)=\varphi^{\prime}(\psi(t))\,\psi^{\prime}(t)=f_{N}(y_{\varphi\circ\psi(t)})\,f_{N}(z_{\psi(t)})^{-1}=1,\qquad\text{for a.e. }t\in(0,T).$ (52)
Recalling that $\lim_{s\downarrow 0}\varphi(s)=0$ and $\lim_{t\downarrow 0}\psi(t)=0$, we deduce that $\varphi\circ\psi(t)=t$ for every $t\in(0,T)$ and, as a consequence, we conclude that $\mathcal{R}_{2}\big{(}\mathcal{R}_{1}\big{(}(y_{t})_{t\in(0,T)}\big{)}\big{)}=(y_{t})_{t\in(0,T)}$. Therefore, we have shown that $\mathcal{R}_{1}\circ\mathcal{R}_{2}=\operatorname{id}_{\mathcal{C}^{\prime}}$, analogously we can prove that $\mathcal{R}_{2}\circ\mathcal{R}_{1}=\operatorname{id}_{\mathcal{C}^{\prime\prime}_{N}}$. ∎
@@ 3.2  Local characterization and self-improvement of $\operatorname{EVI}_{K,N}$ gradient curves
Proof.
Directly from the definition of directional derivative of $f_{N}$ and the $(K,N)$-convexity of $f$, we obtain the following inequality
$f^{\prime}_{N}(x_{0};x)\leq\liminf_{t\downarrow 0}\dfrac{\sigma_{K,N}^{(t)}(\mathsf{d})f_{N}(x_{1})+\big{(}\sigma_{K,N}^{(1-t)}(\mathsf{d})-1\big{)}f_{N}(x_{0})}{t}.$
Hence, the validity of the following identities
$\lim_{t\downarrow 0}\dfrac{\sigma_{K,N}^{(t)}(\mathsf{d})}{t}=\dfrac{\mathsf{d}}{\mathfrak{s}_{K,N}(\mathsf{d})}\quad\text{ and }\quad\lim_{t\downarrow 0}\dfrac{\sigma_{K,N}^{(1-t)}(\mathsf{d})-1}{t}=-\dfrac{\mathfrak{c}_{K,N}(\mathsf{d})}{\mathfrak{s}_{K,N}(\mathsf{d})}\mathsf{d}$
leads to the desired conclusion. ∎
@@
where the last equality is a consequence of Lemma 3.5.
@@ 3. $(iii)$
    where $\tilde{M}\coloneqq-\frac{K}{N}e^{-M/N}$ and $M:=\max\{f(z_{r}),f(\tilde{z}_{r})\}$.
@@
$\mathfrak{s}_{K,N}(\theta/2)^{2}=-\frac{N}{2K}\bigg{(}\mathfrak{c}_{K,N}(\theta)-1\bigg{)},$ (79)
@@
$\begin{split}\sigma_{K,N}^{(1-s)}(\mathsf{d})f_{N}(x_{0})&+\sigma_{K,N}^{(s)}(\mathsf{d})f_{N}(x_{1})\\ &\geq\frac{f_{N}(\gamma_{s}^{t})}{(e^{Kt}-1)}\bigg{[}\sigma_{K,N}^{(1-s)}(\mathsf{d})\big{(}e^{Kt}\mathfrak{c}_{K,N}(\mathsf{d}(\gamma_{s}^{t},x_{0}))-\mathfrak{c}_{K,N}(s\mathsf{d})\big{)}\\ &\qquad\qquad\qquad\qquad+\sigma_{K,N}^{(s)}(\mathsf{d})\big{(}e^{Kt}\mathfrak{c}_{K,N}(\mathsf{d}(\gamma_{s}^{t},x_{1}))-\mathfrak{c}_{K,N}((1-s)\mathsf{d})\big{)}\bigg{]},\end{split}$ (80)
for every $t$ sufficiently small. Now, as a consequence of the trigonometric sum formulas, we observe that
$\sigma_{K,N}^{(1-s)}(\mathsf{d})\mathfrak{c}_{K,N}(s\mathsf{d})+\sigma_{K,N}^{(s)}(\mathsf{d})\mathfrak{c}_{K,N}((1-s)\mathsf{d})=1.$ (81)
Moreover, we take adventage of the following trigonometric inequalities stated at the end of the proof of [EKS15, Thm. 2.23]: for every $\alpha,\alpha^{\prime}\geq 0$ and $\varepsilon,\varepsilon^{\prime}\in[-\pi,\pi]$ such that $\varepsilon+\varepsilon^{\prime}\geq 0$, setting $\beta=\alpha+\varepsilon$ and $\beta^{\prime}=\alpha^{\prime}+\varepsilon^{\prime}$, we have that
$\begin{split}\sin(\alpha)\cos(\beta^{\prime})+\cos(\beta)\sin(\alpha^{\prime})&\leq\sin(\alpha+\alpha^{\prime}),\\ \sinh(\alpha)\cosh(\beta^{\prime})+\cosh(\beta)\sinh(\alpha^{\prime})&\geq\sinh(\alpha+\alpha^{\prime}).\end{split}$ (82)
Applying these with $\alpha=(1-s)\mathsf{d}$, $\alpha^{\prime}=s\mathsf{d}$, $\varepsilon=\mathsf{d}(\gamma_{s}^{t},x_{1})-(1-s)\mathsf{d}$ and $\varepsilon^{\prime}=\mathsf{d}(\gamma_{s}^{t},x_{0})-s\mathsf{d}$, in which case $\varepsilon+\varepsilon^{\prime}\geq 0$ by the triangular inequality, we deduce that
$\sigma_{K,N}^{(1-s)}(\mathsf{d})\mathfrak{c}_{K,N}(\mathsf{d}(\gamma_{s}^{t},x_{0}))+\sigma_{K,N}^{(s)}(\mathsf{d})\mathfrak{c}_{K,N}(\mathsf{d}(\gamma_{s}^{t},x_{1}))\,\begin{cases}\leq 1&\text{if }K<0,\\ \geq 1&\text{if }K>0.\end{cases}$ (83)
@@ 5  $\operatorname{EVI}$ and $\operatorname{EDI}$
$f(y_{t})\leq\lim_{\epsilon\downarrow 0}f(y_{\epsilon})-\int_{0}^{t}\left[\frac{1}{2}|\dot{y}_{s}|^{2}+\frac{1}{2}|D^{-}f|^{2}(y_{s})\right]\,\mathrm{d}s\qquad\text{for every }t\in(0,T).$ (EDI)
@@ 3. $(c)$
    $-\frac{\,\mathrm{d}}{\,\mathrm{d}t}f(y_{t})\geq\frac{1}{2}|\dot{y}_{t}|^{2}+\frac{1}{2}|D^{-}f|^{2}(y_{t})\qquad\text{for }\mathscr{L}^{1}\text{-a.e. }t\in(0,T).$ (85)
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}f(y_{t})=-|D^{-}f|^{2}(y_{t})=-|\dot{y}_{t}|^{2}\qquad\text{for }\mathscr{L}^{1}\text{-a.e. }t\in(0,T).$ (86)
@@
$f(y_{t})=\lim_{\epsilon\downarrow 0}f(y_{\epsilon})-\int_{0}^{t}\left[\frac{1}{2}|\dot{y}_{s}|^{2}+\frac{1}{2}|D^{-}f|^{2}(y_{s})\right]\,\mathrm{d}s\qquad\text{for every }0<t<T.$ (EDE)
@@
Fix any $y\in\mathcal{D}[f]$ and $R>0$. We start by observing that, for every $z$ in a neighborhood of $y$,
@@
$-\frac{N}{f_{N}(y)}\frac{f_{N}(\gamma_{t})-f_{N}(y)}{\mathsf{d}(y,\gamma_{t})}\leq$ $-\frac{N}{f_{N}(y)}\frac{\sigma_{K,N}^{(t)}(\mathsf{d})f_{N}(z)-(1-\sigma_{K,N}^{(1-t)}(\mathsf{d}))f_{N}(y)}{t\mathsf{d}}.$ (95)
Letting $t\to 0$ in the right-hand side above, we obtain
$\lim_{t\downarrow 0}-\frac{N}{f_{N}(y)}\frac{\sigma_{K,N}^{(t)}(\mathsf{d})f_{N}(z)-(1-\sigma_{K,N}^{(1-t)}(\mathsf{d}))f_{N}(y)}{t\mathsf{d}}=$ $-\frac{N}{f_{N}(y)}\bigg{[}\frac{1}{\mathfrak{s}_{K,N}(\mathsf{d})}f_{N}(z)-\frac{\mathfrak{c}_{K,N}(\mathsf{d})}{\mathfrak{s}_{K,N}(\mathsf{d})}f_{N}(y)\bigg{]}$
$=$ $-\frac{N}{f_{N}(y)}\bigg{[}\frac{f_{N}(z)-f_{N}(y)}{\mathfrak{s}_{K,N}(\mathsf{d})}-\frac{1-\mathfrak{c}_{K,N}(\mathsf{d})}{\mathfrak{s}_{K,N}(\mathsf{d})}f_{N}(y)\bigg{]}$
$=$ $\frac{N}{\mathfrak{s}_{K,N}(\mathsf{d})}\left[1-\frac{f_{N}(z)}{f_{N}(y)}\right]-K\,\frac{\mathfrak{s}_{K,N}(\mathsf{d}/2)}{\mathfrak{c}_{K,N}(\mathsf{d}/2)}f_{N}(y),$
@@
$\mathfrak{l}^{\Gamma}_{f}(y):=\sup_{z\in\Gamma\setminus\{y\}}\frac{\big{(}f(z)-f(y)\big{)}^{-}}{\mathsf{d}(z,y)}.$
@@
$\mathfrak{l}^{\Gamma}_{f}(y)\leq|D^{-}f|(y)+|K|\max\bigg{\{}\frac{\mathfrak{s}_{K,N}(R/2)}{\mathfrak{c}_{K,N}(R/2)},\frac{\mathfrak{s}_{-K,N}(R/2)}{\mathfrak{c}_{-K,N}(R/2)}\bigg{\}},\qquad\forall y\in\Gamma.$
@@
Since $\left(y_{t}\right)_{t\in(0,T)}$ is of maximal slope, by (85) we have $\frac{\,\mathrm{d}}{\,\mathrm{d}t}f(y_{t})\leq 0$ $\mathscr{L}^{1}$-a.e., thus $t\mapsto f(y_{t})$ is non-increasing in light of Definition 5.2b. We deduce that $\left(y_{t}\right)_{t\in(0,T)}\in\mathcal{C}^{\prime}$ by definition (40) of $\mathcal{C}^{\prime}$.
Consider the reparametrizations $\mathcal{R}_{1}\big{(}\left(y_{t}\right)_{t\in(0,T)}\big{)}$ and $\mathcal{R}_{1}\big{(}\left(y^{\prime}_{t}\right)_{t\in(0,T)}\big{)}$. Since $\mathcal{R}_{1}$ is injective on $\mathcal{C}^{\prime}$ by Proposition 3.1, it suffices to show that $\mathcal{R}_{1}\big{(}\left(y_{t}\right)_{t\in(0,T)}\big{)}=\mathcal{R}_{1}\big{(}\left(y^{\prime}_{t}\right)_{t\in(0,T)}\big{)}$.
@@
$-\frac{\,\mathrm{d}}{\,\mathrm{d}t}f_{N}(z_{t})$ $=-\frac{f_{N}(z_{t})}{N}\left[-\frac{\,\mathrm{d}}{\,\mathrm{d}t}f(z_{t})\right]=\varphi^{\prime}(t)\left[-\frac{\,\mathrm{d}}{\,\mathrm{d}t}f(z_{t})\right]$ (97)
@@
$-\frac{\,\mathrm{d}}{\,\mathrm{d}t}f(z_{t})=-\frac{\,\mathrm{d}}{\,\mathrm{d}t}f(y_{\varphi(t)})=\varphi^{\prime}(t)\frac{\,\mathrm{d}}{\,\mathrm{d}s}f(y_{s})\big{\rvert}_{s=\varphi(t)}\geq\varphi^{\prime}(t)\left[\frac{1}{2}\left\lvert\dot{y}_{\varphi(t)}\right\rvert^{2}+\frac{1}{2}\left\lvert D^{-}f\right\rvert^{2}(y_{\varphi(t)})\right].$ (98)
@@
$-\frac{\,\mathrm{d}}{\,\mathrm{d}t}f_{N}(z_{t})\geq\varphi^{\prime}(t)^{2}\left[\frac{1}{2}\left\lvert\dot{y}_{\varphi(t)}\right\rvert^{2}+\frac{1}{2}\left\lvert D^{-}f\right\rvert^{2}(y_{\varphi(t)})\right]=\frac{1}{2}\left\lvert\dot{z}_{t}\right\rvert^{2}+\frac{1}{2}\left\lvert D^{-}f_{N}\right\rvert^{2}(z_{t}).$ (99)
@@
Conflict Of Interest and Data Availability
The authors certify that they have NO affiliations with or involvement in any organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript. Moreover, this manuscript has no associated data.